the bottom of pool. Each finds Four people are allowed to pick pennies from Y, dollars worth of pennies during the X, minutes the four people are allowed in the pool: 1.7 a Minutes, X Dollars Found, Y, 2.00 5.00 6.00 7.00 Person (a) What is the centroid for this sample? (b) What is b, of the least squares line Y bo + bX? (c) What is bo of the least squares line Y bo t bX? (d) Interpretation: People tend to find aboutfor each additional minute in the pool.
Answer #1

Ans a)

Centroid for this sample would be the dollars per minute collected therefore

Sum of dollars/total time in minutes=20/12=1.67 dollars per minute

Ans B

Mean of Y=20/4=5

Mean of X=12/4=3

y=Y-Mean(Y) x=Mean(X)

-3 -2

0 -1

1 0

2. 3

Sum(xy)=12

Sum(x^2)=14

Beta1=12/14=5/7

Beta0=Mean(Y)-Beta1*Mean(X)=5-(5/7)*3=20/7

For each additional minute

We have 5/7 increment in dollar